LASER SPECTROSCOPY OF NICKEL CYANIDE, NiCN, IN THE VISIBLE REGION

\underline{CHRISTOPHER T. KINGSTON}, ANTHONY J. MERER, Department of Chemistry, University of British Columbia, 2036 Main Mall, Vancouver, BC, Canada, V6T 1Z1; THOMAS D. VARBERG, Department of Chemistry, Macalester College, St. Paul, MN 55105.

The electronic spectrum of NiCN has been observed in the visible region following the reaction of laser-ablated nickel with either (CN)_{2} or CH_{3}CN under supersonic jet-cooled conditions. High resolution LIF experiments have shown that the ground state is ^{2}Δ_{5/2}, with B_{0} = 0.1444334(30) cm^{−1}; using data taken with CH_{3}C^{15}N, the bond lengths are found to be r_{0} (Ni-C) = 1.8292(37) Å and r_{0} (C-N) = 1.1591(52) Å. The vibrational structure of the ground state is an excellent example of strong Fermi resonance. Four excited electronic states have been identified. They are assigned as Ã^{2}Δ_{5/2} at 16607 cm^{−1}, B̃^{2}Π_{3/2} at 16686 cm^{−1}, C̃^{2}Φ_{7/2} at 17813 cm^{−1} and D̃^{2}Φ_{5/2} at 19129 cm^{−1}. The electronic structure of NiCN turns out to be remarkably similar to that of NiH ^{a}.

aS.A. Kadavathu, R. Scullman, R.W. Field, J.A. Gray and M. Li, J. Mol. Spectrosc. \underline{147}, 448 (1991)